# MEM1: Learning to Synergize Memory and Reasoning for Efficient Long-Horizon Agents

### v1

@@ A.4.1 Metrics
In QA tasks, we use exact match (EM) as both the verifiable reward for the RL pipeline and the evaluation metric for the final output. The final response is extracted from between <answer> and </answer>. In multi-objective settings, the response should contain answers to each question separated by semicolons. If the XML tags are mismatched, or if the number of provided answers does not correspond to the number of questions, a score of is assigned. Otherwise, point is credited for each correct answer. During RL training, we do not provide any other intermediate rewards or format penalties, as we find that such manual interventions can interfere with the agent’s learning process (see more in Sec.˜4.4).
@@ A.5 Algorithm
We provide an outline of the rollout of MEM1, which actively manages its context in Alg.˜1. Parts of the pseudo-code follow [22].
@@ Appendix C Training Trajectory Analysis of MEM1
We present the training dynamics of the 2-objective QA-trained MEM1 in Fig.˜6, where several distinct phases emerge during the learning process. In the initial exploration phase (first 50 steps), the agent demonstrates little task proficiency. The reward remains consistently low, while the entropy loss is high, suggesting random or undirected behavior. The ratio of valid actions hovers around 0.55, indicating that the agent frequently fails to follow the expected output format. During this period, MEM1 has not yet learned to reliably use the required structure involving <query> and <answer> tags.
@@ D.1 RL Generalizes Better Than SFT
A natural question arises: can Supervised Fine-Tuning (SFT) with high-quality trajectories match the performance of reinforcement learning (RL)? To investigate this, we compare MEM1-QA trained via RL against MEM1-QA (SFT), where both models are trained on the 2-objective QA task. Additionally, the SFT model is further trained on 1-objective and 3-objective QA tasks to enhance its generalization ability. As shown in Fig.˜7, the SFT model consistently underperforms compared to its RL counterpart across tasks with varying numbers of questions (objectives). Notably, when the number of objectives exceeds six, the performance of the SFT model collapses, whereas the RL-trained model continues to demonstrate strong robustness and scalability.
@@ D.2 Format Reward Accelerates Convergence but Degrades Final Performance
As shown in Fig.˜7, using the format reward leads to faster convergence during training but results in worse final performance. The format-constrained agent achieves an exact match score of 0.466, compared to 0.709 for MEM1 trained with only outcome-based reward on the same testing set for the 2-objective QA task. Additionally, the format-constrained agent generates fewer tokens, with an average peak of 514.9 tokens, whereas the outcome-reward-trained MEM1 reaches an average peak of 640 tokens.

### v2

@@ A.4.1 Metrics
In QA tasks, we use exact match (EM) as both the verifiable reward for the RL pipeline and the evaluation metric for the final output. The final response is extracted from between <answer> and </answer>. In multi-objective settings, the response should contain answers to each question separated by semicolons. If the XML tags are mismatched, or if the number of provided answers does not correspond to the number of questions, a score of is assigned. Otherwise, point is credited for each correct answer. During RL training, we do not provide any other intermediate rewards or format penalties, as we find that such manual interventions can interfere with the agent’s learning process (see more in Sec. 4.4).
@@ A.5 Algorithm
We provide an outline of the rollout of MEM1, which actively manages its context in Alg. 1. Parts of the pseudo-code follow [24].
@@ A.6 MEM1 on Webshop Training Details
A.7 Additional Discussion on the Attention Matrix Design.
We wish to note that our modification to the attention matrix does not fully recover the attention of the original trajectories because of the change in position ids. Specifically, prior works [28, 10, 8] that utilized the attention matrix to compress multiple trajectories mainly targeted tree-exploration, i.e., generating multiple sequences with the same prefix. For these works, on top of the attention matrix, they adjusted the position ids as well, so each trajectory follows a consecutive increasing position ids. However, in MEM1, the prefix does not remain the same because of memory consolidation. This results in each <IS> having two possible position ids, one for the previous turn and one for the next turn. To completely recover the original attention, we need to duplicate each <IS> and assign different position ids to the two copies. However, such duplication can significantly slow down training because the training trajectories are now much longer.
As such, for training efficiency, we do not duplicate the <IS> and assign the position ids for the previous trajectory to each <IS>. While this modification slightly deviates from the “ideal” implementation, effectively, it can be viewed as simply adding white spaces in the training trajectories and has no significant impact on the experimental results.
@@ Appendix C Training Trajectory Analysis of MEM1
We present the training dynamics of the 2-objective QA-trained MEM1 in Fig. 6, where several distinct phases emerge during the learning process. In the initial exploration phase (first 50 steps), the agent demonstrates little task proficiency. The reward remains consistently low, while the entropy loss is high, suggesting random or undirected behavior. The ratio of valid actions hovers around 0.55, indicating that the agent frequently fails to follow the expected output format. During this period, MEM1 has not yet learned to reliably use the required structure involving <query> and <answer> tags.
@@ D.1 RL Generalizes Better Than SFT
A natural question arises: can Supervised Fine-Tuning (SFT) with high-quality trajectories match the performance of reinforcement learning (RL)? To investigate this, we compare MEM1-QA trained via RL against MEM1-QA (SFT), where both models are trained on the 2-objective QA task. Additionally, the SFT model is further trained on 1-objective and 3-objective QA tasks to enhance its generalization ability. As shown in Fig. 7, the SFT model consistently underperforms compared to its RL counterpart across tasks with varying numbers of questions (objectives). Notably, when the number of objectives exceeds six, the performance of the SFT model collapses, whereas the RL-trained model continues to demonstrate strong robustness and scalability.
@@ D.2 Format Reward Accelerates Convergence but Degrades Final Performance
As shown in Fig. 7, using the format reward leads to faster convergence during training but results in worse final performance. The format-constrained agent achieves an exact match score of 0.466, compared to 0.709 for MEM1 trained with only outcome-based reward on the same testing set for the 2-objective QA task. Additionally, the format-constrained agent generates fewer tokens, with an average peak of 514.9 tokens, whereas the outcome-reward-trained MEM1 reaches an average peak of 640 tokens.